# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 5.3. Duals of grid patches
Up to translation, there are $24$ Wang patches among all the Penrose tilings $x\in\mathsf{X}$. We denote these (proto-)Wang patches by $\{w^{a}:a\in{\mathcal{A}}\}$ where ${\mathcal{A}}=\{0,\dots,23\}$.
@@
By Lemma 4.4 it suffices to consider $g_{\boldsymbol{0}}({\boldsymbol{u}})^{*}$ with ${\boldsymbol{u}}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$. These points ${\boldsymbol{u}}$ correspond to the $2$-torus $\mathbb{T}^{2}=\{(u_{2},u_{4})\in[0,1]^{2}\}/\{(u_{2},0)\sim(u_{2},1),(0,u_{4})\sim(1,u_{1})\}$. By Lemma 5.8, we count inequivalent patches under the additional assumption that $y({\boldsymbol{u}})\in Y_{1}$. Figure 10 shows a bifurcation diagram, (called a Markov partition in [16]), consisting of a division of $\mathbb{T}^{2}$ into $24$ cells. A point $(u_{2},u_{4})$ is in the interior of a cell if $g_{{\boldsymbol{0}}}({\boldsymbol{u}})$ has only $2$-fold crossings. When $(u_{2},u_{4})$ on an edge or vertex, $g_{{\boldsymbol{0}}}({\boldsymbol{u}})$ has $3$-fold or $5$-fold crossings. The cells are found by solving systems of linear equations. ∎
The $24$ (proto-)Wang patches, up to translation, are shown in Figure 9. Pictures of these patches appeared in the first author’s dissertation [17], and also in [16].
Figure 10. The bifurcation diagram showing the parameter values $(u_{2},u_{4})\in\mathbb{T}^{2}\sim[0,1]^{2}$ corresponding the 24 Wang patches. Each cell is labeled with $a\in{\mathcal{A}}$ for the corresponding Wang patch $w^{a}$ (Figure 9). The vertices along each edges of the square are at 00, $2-\gamma\approx 0.381967$, $\gamma-1\approx 0.618034$, $4-2\gamma\approx 0.763932$, and $1$, where $\gamma=(1/2)(1+\sqrt{5})$. The courser division into four cells by the lines $u_{2}=2-\gamma$ and $u_{4}=2-\gamma$ corresponds to the values $0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1$ of the symbol $\sigma(y({\boldsymbol{u}}))$ where ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$ with $u_{2}=\{-u_{2}-u_{3}\}$ (read left to right, bottom to top). In other words, this four cell partition defines the $\mathbb{Z}^{2}$ Sturmian shift $S\otimes S$.
@@
Recall (Lemma 5.5) that for $x\in\mathsf{X}$, the points ${\boldsymbol{d}}_{\boldsymbol{n}}$ for ${\boldsymbol{n}}\in\mathbb{Z}^{2}$ are the upper right vertices of the type ${\boldsymbol{v}}^{\prime}_{0}\wedge{\boldsymbol{v}}^{\prime}_{1}$ corner tiles $D_{\boldsymbol{n}}$ of the Wang patches $w_{\boldsymbol{n}}(x)$. For each $n_{0}\in\mathbb{Z}$, we define the $n_{0}$th horizontal edge curve $\eta_{n_{0}}(x)$ to be the piecewise linear curve connecting the vertices $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{1}\in\mathbb{Z}}$. Similarly, for $n_{1}\in\mathbb{Z}$ we define the $n_{1}$th vertical edge curve $\eta_{n_{1}}(x)$ to be the piecewise linear curve connecting the vertices $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{0}\in\mathbb{Z}}$.
The edge curves $\{\eta_{n_{0}}(x)\}_{n_{0}\in\mathbb{Z}}$ and $\{\eta_{n_{1}}(x)\}_{n\in\mathbb{Z}}$ partition $\mathbb{R}^{2}$ into a tiling by tiles by tetragons (i.e., quadrilaterals):
@@
For $x\in\mathsf{X}$ and ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, let $U_{\boldsymbol{n}}(x)$ be the tile bounded by the tetragon with vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, and ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, (the edges of $U_{\boldsymbol{n}}(x)$ are segments of the edge curves $\eta_{n_{0}}(x)$, $\eta_{n_{0}+1}(x)$, $\eta_{n_{1}}(x)$ and $\eta_{n_{1}+1}(x)$ in $x$). We call $U_{\boldsymbol{n}}$ an unmarked tetragon Penrose tile, and call $u=\{U_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ an unmarked tetragon Penrose tiling.
There is a unique tetragon tile $U_{\boldsymbol{n}}(x)$ in $u$ corresponding to each Wang patch $w_{\boldsymbol{n}}(x)$ in $x$ (see Figure 11). In Section 6 we will show that there are, up to translation, $11$ different tetragon prototiles $\{U^{t}:t=0,\dots,10\}$, so some of the tetragon tiles correspond to more than one Wang patches $w^{a}$. We will return to these tilings in Section 6.
Figure 11. Swatch of Penrose tiling $x\in\mathsf{X}$ (a) with type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tiles $D_{\boldsymbol{n}}$ shaded to show corners of Wang patches $w^{\boldsymbol{n}}(x)$. (b) Edge curves connect the upper right vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$ of $D_{\boldsymbol{n}}$ delineating unmarked tetrgon Wang tiles $U_{\boldsymbol{n}}$.
@@
Suppose $x=y({\boldsymbol{u}})^{*}\in\mathsf{X}_{1}$ is a nonsingular Penrose tiling. For $n_{0},n_{1}\in\mathbb{Z}$ let $\ell_{0,n_{1}}(u_{0})$ and $\ell_{1,n_{0}}(u_{1})$ be the corresponding 00 and $1$-grid lines in in $y({\boldsymbol{u}})$, and let $\eta_{n_{1}}(x)$ and $\eta_{n_{0}}(x)$ be the vertical and horizontal edge curve through $x$. Then $||\ell_{0,n_{1}}(u_{0})-\eta_{n_{1}}(x)||,||\ell_{1,n_{0}}(u_{1})-\eta_{n_{0}}(x)||<2/3$ (in the sense of the Euclidean uniform closeness of the two curves). The same results hold for singular Penrose tilings $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, if $y({\boldsymbol{u}})=\varphi(x)$.

### v2

@@ 5.3. Duals of grid patches
Up to translation, there are $24$ Wang (proto-) patches among all the Penrose tilings $x\in\mathsf{X}$.
We denote these Wang (proto-) patches by $\{w^{a}:a\in{\mathcal{A}}\}$ where ${\mathcal{A}}=\{0,\dots,23\}$. These $24$ Wang patches are shown in Figure 9. Pictures of these patches appeared in the first author’s dissertation [17], and also in [16].
Figure 10. The bifurcation diagram, showing the parameter values $(u_{2},u_{4})\in\mathbb{T}^{2}\sim[0,1]^{2}$ corresponding the 24 Wang patches. Each cell is labeled with $a\in{\mathcal{A}}$ for the corresponding Wang patch $w^{a}$ (Figure 9). The vertices along each edges of the square are at 00, $2-\gamma\approx 0.381967$, $\gamma-1\approx 0.618034$, $4-2\gamma\approx 0.763932$, and $1$. The coarser division of $[0,1]^{2}$ into four cells by the lines $u_{2}=2-\gamma$ and $u_{4}=2-\gamma$ corresponds to the values $0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1$ of the symbol $\sigma(y({\boldsymbol{u}}))$ where ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$ with $u_{2}+u_{3}+u_{4}=0$, corresponding to the $\mathbb{Z}^{2}$ Sturmian shift $S\otimes S$.
Proof of Proposition 5.9.
By Lemma 5.8, all the Wang patches $w^{a}$, up to translation, can be obtained as duals $g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ of all the grid patches $g_{\boldsymbol{n}}({\boldsymbol{u}})$ that occur in all $y({\boldsymbol{u}})\in Y_{1}$. This means all crossings in $y({\boldsymbol{u}})$ are 2-fold, and in particular, this holds for any grid patch in $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})$. We will call a grid patch that has only 2-fold crossings a nonsingular grid patch. Now, if $g_{\boldsymbol{n}}({\boldsymbol{u}})$ is a nonsingular grid patch in $y({\boldsymbol{u}})\in Y$, (i.e., $y({\boldsymbol{u}})$ is not necessarily nonsingular), then since $Y_{1}$ is dense $G_{\delta}$, there exists a small perturbation of ${\boldsymbol{u}}$, which does not change $g_{\boldsymbol{n}}({\boldsymbol{u}})$ enough to change its dual $g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$, so that with this new ${\boldsymbol{u}}$, we have $y({\boldsymbol{u}})\in Y_{1}$.
Given any grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})\in Y$ we can translate it to the origin, using Lemma 4.6 to adjust ${\boldsymbol{u}}$, so that for this new ${\boldsymbol{u}}$, the patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ is translated to the origin and becomes $g_{\boldsymbol{0}}({\boldsymbol{u}})$, with its support satisfying $R_{\boldsymbol{0}}=R$, where $R$ is as in (5.2). The new value of ${\boldsymbol{u}}$ thus has the form ${\boldsymbol{u}}=(0,0,u_{2},u_{3},u_{4})$, where we assume without loss of generality that $(u_{2},u_{3})\in[0,1]^{2}$ with $u_{2}+u_{3}+u_{4}=0$. Thus we think of $g_{\boldsymbol{0}}({\boldsymbol{u}})$ as depending on $(u_{2},u_{4})\in[0,1]^{2}$, but we also observe that $g_{\boldsymbol{0}}({\boldsymbol{u}})$ is the same for $(u_{2},0)$ and $(u_{2},1)$, and for $(0,u_{4})$ and $(1,u_{4})$, so we can also regard $(u_{2},u_{4})\in\mathbb{T}^{2}$.
Figure 10 shows a partition of $[0,1]^{2}$ into $24$ cells, $46$ edges, and $23$ vertices. We call this the bifurcation diagram for Wang patches, noting that the same diagram, called a Markov partition for Wang tiles, was obtained in [16]. The cells in the bifurcation diagram are defined by the condition that a point $(u_{2},u_{4})$ is in the interior of a cell if and only if the corresponding grid patch $g_{{\boldsymbol{0}}}({\boldsymbol{u}})$ is nonsingular. It follows that the dual $g_{{\boldsymbol{0}}}({\boldsymbol{u}})^{*}$ is a constant Wang patch $w^{a}$ for all $(u_{2},u_{4})$ in the interior of each cell. By the discussion in the previous two paragraphs, the duals $g_{{\boldsymbol{0}}}({\boldsymbol{u}})^{*}$ range over all possible Wang patches $w^{a}$ up to translation. We can see each of these Wang patches by drawing the dual for the grid patch $g_{\boldsymbol{0}}({\boldsymbol{u}}_{0})$ corresponding to a single point ${\boldsymbol{u}}_{0}$ in each cell. We find that each cell results in a distinct Wang patch. Thus there are 24 distinct Wang patches, which are the Wang patches $w^{a}$, $a\in{\mathcal{A}}$, shown in Figure 9. The numbering $a\in{\mathcal{A}}$ corresponds to the labels of the cells in Figure 10. In particular, these are the 24 Wang patches described in [11].
It remains to show that the bifurcation diagram is correct. Figure 10 shows five vertices along each side of $[0,1]^{2}$ located at $\eta_{0}=0$, $\eta_{1}:=2-\gamma\approx 0.381967$, $\eta_{2}:=\gamma-1\approx 0.618034$, $\eta_{3}:=4-2\gamma=3-\sqrt{5}\approx 0.763932$, and $\eta_{4}=1$. We will now locate all the possible multiple crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$. The remainder of the proof is organized in five steps.
Step 1. It is easy to see that there are $5$-fold crossings $g_{\boldsymbol{0}}({\boldsymbol{u}})$ exactly when $(u_{2},u_{4})$ in $[0,1]^{2}$ is one of the nine vertices corresponding to the partition of $[0,1]^{2}$ into four cells by the vertical line $u_{2}=\eta_{1}$ and the horizontal line $u_{4}=\eta_{1}$, (in $\mathbb{T}^{2}$ these identify to four vertices). In subsequent discussions we implicitly ignore these nine points, and look only for $3$-fold crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$.
Altogether, there are ten possible types of $3$-fold crossings, which we denote $012$, (meaning a crossing of a 00-, $1$- and a $2$-grid line), $013$, $014$, $023$, $024$, $034$ $123$, $124$, $134$, and $234$.
Step 2. It is easy to see that if $(u_{2},u_{4})$ lies along the left or right side of $[0,1]^{2}$ (on the line $u_{2}=0$ or $u_{2}=1$) then there are $2$-grid lines through the lower left and lower right corners of $R_{\boldsymbol{0}}$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, as shown in Figure 8(c). Similarly, for points along the vertical line $u_{2}=\eta_{1}$, there are $2$-grid lines through the upper left and upper right corners of $R_{\boldsymbol{0}}$. These account for all possible $3$-fold crossings type of $012$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$. In a similar way, for $(u_{2},u_{4})$ along any of the three horizontal lines $u_{2}=0$, $u_{2}=\eta_{1}$ or $u_{2}=1$, there are $3$-fold crossings of type $014$ at corners of $R_{\boldsymbol{0}}$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, and there can be no other crossings of this type.
Step 3. For all $(u_{2},u_{4})$ along the five lines of slope $-1$ in the bifurcation diagram, (i.e., the lines $u_{2}+u_{4}=\eta_{1}$, $u_{2}+u_{4}=\eta_{2}$, $u_{2}+u_{4}=1$, $u_{2}+u_{4}=2\eta_{2}$ and $u_{2}+u_{4}=2\eta_{3}$) there are $3$-fold crossings type $013$ at corners of $R_{\boldsymbol{0}}$. For two of these, $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$, the $3$-grid lines cross the upper left and lower right corners. For the other three, the $3$-grid lines cross the lower left and upper right corners of $R_{\boldsymbol{0}}$. No other crossings of type $013$ are possible.
Step 4. We show that for $(u_{2},u_{4})\in[0,1]^{2}$, all possible type $234$ crossings occur along (two of the) slope $-1$ lines from Step 3: $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$. By Section 4.5, Case A, all $3$-fold crossings in $y({\boldsymbol{u}})$ occur along spines with both wide and narrow $3$-fold crossings (Figure 6(a)). In Step 3 above, the spine is a $3$-grid line. The type $013$ in Step 3 are wide crossings, so the narrow crossings are type $234$. These occur in $g_{\boldsymbol{0}}({\boldsymbol{u}})$ in the cases where the $3$-grid line goes through the lower left and upper right corners of $R_{\boldsymbol{0}}$: $u_{2}+u_{4}=\eta_{1}$ and $u_{2}+u_{4}=2\eta_{2}$.
Step 5. Finally, we discuss the six types of $3$-fold crossings that have the form $0ij$ or $1ij$ for $i,j=2,3,4$, $i<j$. Any such crossing occurs along an edge of $R_{\boldsymbol{0}}$. By (4.1), a grid line $\ell_{i,k}(u_{i})$ has the equation ${\boldsymbol{v}}_{i}\cdot(x,y)=(-u_{i}+k)$. So if $\ell_{i,k}(u_{i})\cap\ell_{j,k^{\prime}}(u_{j})=\{(x,y)\}$ then $V_{i,j}(x,y)=(-u_{i}+k,-u_{j}+k^{\prime})$, where $V_{i,j}$ is the $2\times 2$ matrix with rows ${\boldsymbol{v}}_{i}$ and ${\boldsymbol{v}}_{j}$. Recalling that we identify $(x,y)$ with $[x,y]^{t}$, we define an affine map $A_{i,j,k,k^{\prime}}(x,y):=M_{i,j}V_{i,j}(x,y)-(k,k^{\prime})$, where
$M_{2,4}:=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right),\ \ \ M_{2,3}:=\left(\begin{matrix}1&0\\ -1&-1\end{matrix}\right),\ \ \ M_{3,4}:=\left(\begin{matrix}-1&-1\\ 0&1\end{matrix}\right).$ (5.4)
Then $A_{i,j,k,k^{\prime}}$ maps a crossing $(x,y)$ of $\ell_{i,k}(u_{i})$ and $\ell_{j,k^{\prime}}(u_{j})$ ($i,j=2,3,4$, $i<j$) to a point $(u_{2},u_{4})$ in the plane of the bifurcation diagram. The matrices $M_{i,j}$ correspond to the fact that $u_{2}+u_{3}+u_{4}=0$. When the map $A_{i,j,k,k^{\prime}}$ is applied to the boundary $\partial R_{\boldsymbol{0}}$ of $R_{\boldsymbol{0}}$ the image is a parallelogram in the $(u_{2},u_{4})$-plane whose intersections with $[0,1]^{2}$ are the values of $(u_{2},u_{4})$ in the bifurcation diagram that correspond to $3$-fold crossings of type $0ij$ or $1ij$ in $g_{\boldsymbol{0}}({\boldsymbol{u}})$.
For each type $ij\in\{24,23,34\}$, Figure 11 shows $A_{i,j,k,k^{\prime}}(\partial R_{\boldsymbol{0}})$ for each of the values of $(k,k^{\prime})$ so that $A_{i,j,k,k^{\prime}}(\partial R_{\boldsymbol{0}})\cap[0,1]^{2}\neq\emptyset$. The caption describes which sides correspond to type $0ij$ and $1ij$ and which $(k,k^{\prime})$ are used.
We conclude by observing that steps 1 - 5 account for all possible $3$-fold and $5$-fold crossings in $g_{\boldsymbol{0}}({\boldsymbol{u}})$, and also account for all the segments in Figure 10.. This shows that the bifurcation diagram is correct. ∎
Figure 11. Images of $\partial R_{\boldsymbol{0}}$ under $A_{i,j,k,k^{\prime}}$ in the $(u_{2},u_{4})$-plane are parallelograms (black). Only those choices of $k,k^{\prime}$ resulting in images that meet $[0,1]^{2}$ (grey) are shown. (a) Images under $A_{2,4,0,0}$, $A_{2,4,0,1}$, and $A_{2,4,1,0}$. The slope $-\gamma$ sides correspond to type $024$ and slope $-1/\gamma$ sides to type $124$. (b) Images under $A_{2,3,0,-1}$, $A_{2,3,-1,-1}$, and $A_{2,3,0,-2}$. The slope $-\gamma$ sides correspond to type $123$, and slope 00 sides (including part of $\partial[0,1]^{2}$) correspond to type $023$. (c) Images under $A_{3,4,-2,0}$, $A_{3,4,-1,0}$, and $A_{3,4,-1,-1}$. The slope $-1/\gamma$ sides correspond to type $034$ and vertical sides (including part of $\partial[0,1]^{2}$) correspond to type $134$.
Definition 5.10.
Fix $x\in\mathsf{X}$. For each $n_{1}\in\mathbb{Z}$ we define the $n_{1}$th bent 00-grid line in $x$, denoted $\lambda_{0,n_{1}}$, to be the piecewise linear curve connecting the points $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{0}\in\mathbb{Z}}$ (where the points ${\boldsymbol{d}}_{\boldsymbol{n}}$ are as defined in Lemma 5.5). Similarly, for each $n_{0}\in\mathbb{Z}$, we define the $n_{0}$th bent $1$-grid line in $x$, denoted $\lambda_{1,n_{0}}$, to be the piecewise linear curve connecting the vertices $\{{\boldsymbol{d}}_{(n_{0},n_{1})}\}_{n_{1}\in\mathbb{Z}}$.
The next lemma shows that bent 00- and $1$-grid lines $\lambda_{1,n_{0}}$ and $\lambda_{n_{1},0}$ closely follow the corresponding grid lines $\ell_{1,n_{0}}$ and $\ell_{n_{1},0}$.
Lemma 5.11.
Suppose $x\in\mathsf{X}_{1}$ and $x=y({\boldsymbol{u}})^{*}$ for $y({\boldsymbol{u}})\in Y_{1}$. Then $||\ell_{0,n_{1}}-\lambda_{0,n_{1}}||,||\ell_{1,n_{0}}-\lambda_{1,n_{0}}||<2/3$ (in the sense of the Euclidean uniform closeness of the two curves). The same results hold for singular Penrose tilings $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, where $y({\boldsymbol{u}})=\varphi(x)$.
@@
For $x\in\mathsf{X}_{1}$, this follows from Lemma 5.5. For $x\in\mathsf{X}\backslash\mathsf{X}_{1}$, we let $x_{k}=y({\boldsymbol{u}}_{k})^{*}\to x$, where $y({\boldsymbol{u}}_{k})\in Y_{1}$ (see Lemma 4.8) can be chosen so that $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ and $\ell_{0,n_{1}}(u_{0})=\ell_{0,n_{1}}(u^{k}_{0})$ for all $||{\boldsymbol{n}}||$ small. ∎
The bent 00- and $1$-grid lines $\{\lambda_{1,n_{0}}(x)\}_{n_{0}\in\mathbb{Z}}$ and $\{\lambda_{0,n_{1}}(x)\}_{n_{1}\in\mathbb{Z}}$ partition $\mathbb{R}^{2}$ into a tiling by tetragons (i.e., quadrilaterals):
@@
For $x\in\mathsf{X}$ and ${\boldsymbol{n}}=(n_{0},n_{1})\in\mathbb{Z}^{2}$, let $U_{\boldsymbol{n}}(x)$ be the tile bounded by the tetragon with vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}}$, ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{0}+{\boldsymbol{e}}_{1}}$, and ${\boldsymbol{d}}_{{\boldsymbol{n}}+{\boldsymbol{e}}_{1}}$, (the edges of $U_{\boldsymbol{n}}(x)$ are segments of the edge curves $\lambda_{1,n_{0}}$, $\lambda_{1,n_{0}+1}$, $\lambda_{0,n_{1}}$ and $\lambda_{0,n_{1}+1}$ in $x$). We call $U_{\boldsymbol{n}}(x)$ an unmarked tetragon Penrose tile, and call $u(x):=\{U_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ an unmarked tetragon Penrose tiling.
There is a unique tetragon tile $U_{\boldsymbol{n}}(x)$ in $u(x)$ corresponding to each Wang patch $w_{\boldsymbol{n}}(x)$ in $x$ (see Figure 12). In Section 6 we will show that there are, up to translation, $11$ different tetragon prototiles $\{U^{b}:b=0,\dots,10\}$, so some of the tetragon tiles correspond to more than one Wang patche $w^{a}$. We will return to these tilings in Section 6.
Figure 12. Swatch of Penrose tiling $x\in\mathsf{X}$ (a) with type ${\boldsymbol{v}}_{0}^{\prime}\wedge{\boldsymbol{v}}^{\prime}_{1}$ tiles $D_{\boldsymbol{n}}$ shaded to show corners of Wang patches $w^{\boldsymbol{n}}(x)$. (b) Edge curves connect the upper right vertices ${\boldsymbol{d}}_{\boldsymbol{n}}$ of $D_{\boldsymbol{n}}$ delineating unmarked tetragon Wang tiles $U_{\boldsymbol{n}}$.
@@
For each Wang patch $w^{a}$, $a\in{\mathcal{A}}$, we define its symbol $\sigma(w^{j})=z_{0}\otimes z^{\prime}_{0}\in\{0\otimes 0,0\otimes 1,1\otimes 0,1\otimes 1\}$ as follows: If a single direction ${\boldsymbol{v}}_{2}^{\perp}$ trail segment passes through $w^{j}$, we put $z^{\prime}_{0}=0$, but if two distinct direction ${\boldsymbol{v}}_{2}^{\perp}$ trail segments pass through $w^{j}$, we put $z^{\prime}_{0}=1$. Similarly, if a single direction ${\boldsymbol{v}}_{4}^{\perp}$ trail segment passes through $w^{j}$, we put $z_{0}=0$, but if two distinct direction ${\boldsymbol{v}}_{4}^{\perp}$ trail segments pass through $w^{j}$, we put $z_{0}=1$, (compare Definition 5.3).
@@
Let $x\in\mathsf{X}$ be a Penrose tiling covered by Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$. Then $z\otimes z^{\prime}=\{\sigma(w_{{\boldsymbol{n}}}(x))\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ is a $\mathbb{Z}^{2}$ Sturmian sequence. Let $u_{2}=\phi(z)$, $u_{4}=\phi(z^{\prime})$, and let ${\boldsymbol{u}}_{0}=(0,0,u_{2},\{-(u_{2}+u_{4})\},u_{4})$. If $x\in\mathsf{X}_{1}$ then $y({\boldsymbol{u}}_{0})\in Y_{1}$ and $x=T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}}_{0})^{*}$ for some unique ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$. If $x\in\mathsf{X}\backslash\mathsf{X}_{1}$ then $y({\boldsymbol{u}}_{0})^{*}$ has an unfilled worm or unfilled cartwheel, and there is a unique ${\boldsymbol{t}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{t}}_{0}||<4/3$ so that $x$ is $T^{{\boldsymbol{t}}_{0}}y({\boldsymbol{u}})^{*}$ with the worm or cartwheel filled.